Meeting Notes for 6 February 2008

Present: Tom J, Bob., Liz, Mihael, Tom L.;
Absemt: Dan K, Mike W., Randy., Kris, Joao, Eric

Unfortunately Tom's email about the telecon did not include the telephone number. We sent out an email at 11:00 AM with the number, but it may have been too late for the rest of the group to access it. It is the same number every week. We started at 11:00 CST even though we discussed changing it to 11:30 CST because of conflicts the Fermilab team had. Apologies to those who were confused.

Refactored Code – testing

Bob verified that SWIFT shower analysis is now fixed in refactored code. However there are still a couple open problems related to analyses.

  • Saving plots – Mihael thought that he had vixed this, but Bob is still having issues.
  • The need for user-friendly error messages when no events found on shower studies. It should mirror what they should expect to see on the cluster. Mihael noted that this involves the more general problem of propagating errors from scripts back to the user. It needs a more general solution. He needs to think about this.
  • We still need to add some more annotations of the input variables, in particular the choice of VDS/Swift so that users understand the possible problems they might encounter when they choose grid-execution.

Action Item: Liz is going to investigate which input parameters are missing annotations and work out with Mihael how to add these. Tom and Bob will help with composing the annotations.

Mihael has deployed Release Candidate 3 on www12. He is going to try to fix a few of these issues before deploying Release Candidate 4.

Action Item: Bob and Tom will verify that this bug has been fixed in RC4.

Action Item: Milhael will roll out RC4 to www18 to work with the production database for the testers to use.

Action Item: Bob will send a message to e-Lab fellows to test RC4 of the Cosmic e-Lab on this URL after they clear up save plot issue.

Action Item: Liz will make sure all the poster files are there.

Creating a form to harvest help desk ticket information

Tom and Bob still need to create the form. Eric was not available to discuss whether he made any progress on entering this information into a helpdesk system in BOINC.

Proxy Server

We seem to have the problems ironed out. Mihael made an additional change to the timeout, but we should not have this problem any more in the new system.

www13 - deploying refactored code

All of Nick's work has been saved or provided to Mihael. The files that are currently on www13 for Amelia have the text needed to make a refactored version of Amelia. The server does not seem to be running. We would all like to see www13 have the refactored versions.

DAQ Board

Tom L. is visiting Fermilab on Feb. 7th and would like to discuss some issues related the DAQ board.

-- Main.LizQuigg - 07 Feb 2008
Topic revision: r4 - 2008-02-14, LizQuigg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback